# Thermodynamics of ideal spin fluids and pseudo-gauge ambiguity

### v1

@@ Abstract
Conserved currents of relativistic spin fluids derived from microscopic models are known to violate local thermodynamic relations. We present a systematic analysis of pseudo-gauge improvements in ideal spin hydrodynamics and identify a family of pseudo-gauges where standard thermodynamic relations are satisfied. We quantify pseudo-gauge ambiguities in the spin equation of state and derive universal thermodynamic relations that apply to conserved currents in any pseudo-gauge. As an application, we extract the thermodynamic variables and equations of state for free Dirac fermions and scalar fields.
@@
Relativistic heavy-ion collisions produce a short-lived quark-gluon plasma (QGP) that is well described by relativistic hydrodynamics Teaney [2003], Romatschke and Romatschke [2007]. The recent observation of global $\Lambda$-hyperon polarization by the STAR Collaboration demonstrated that the QGP fireball carries substantial angular momentum and exhibits strong vortical structure Adamczyk et al. [2017], STAR Collaboration [2019], ALICE Collaboration [2020, 2022]. This discovery has motivated the development of spin hydrodynamics, a theoretical framework that describes collective spin phenomena and provides novel probes of the QGP’s gradients and rotational dynamics Becattini [2011], Florkowski et al. [2018, 2019a], Montenegro et al. [2017], Hattori et al. [2019], Florkowski et al. [2019b], Becattini and Lisa [2020], Singh et al. [2021], Fukushima and Pu [2021], Hongo et al. [2021], Gallegos et al. [2021], Florkowski et al. [2022], Gallegos et al. [2023], Peng et al. [2021], Cao et al. [2022], Weickgenannt et al. [2022, 2022], Biswas et al. [2023], Florkowski [2025], Huang [2025], Singh [2025], Daher et al. [2025].
@@
Hydrodynamics is a theory of conserved currents. However, these currents admit improvements—the so-called pseudo-gauge transformations—that leave the global conserved charges unchanged Halbwachs [1960], Hehl [1976], Becattini and Tinti [2011], Speranza and Weickgenannt [2021]. As a result, the conserved currents in spin hydrodynamics are not unique, even in equilibrium. In particular, the constitutive relations in generic pseudo-gauges fail to satisfy the standard thermodynamic relations implied by statistical mechanics Becattini et al. [2024], Becattini and Singh [2025], Florkowski and Hontarenko [2025], Drogosz et al. [2024]. This subtlety prevents a direct identification of thermodynamic variables of spin hydrodynamics through naive comparison with the conserved currents obtained from microscopic models or experimental measurements.
@@
For instance, the particle current admits a pseudo-gauge ambiguity $J^{\mu}\to J^{\mu}+\partial_{\sigma}M^{\sigma\mu}_{J}$, for some antisymmetric tensor $M^{\mu\nu}_{J}$. This shifts the particle density in the fluid’s rest frame as $n\to n-u_{\mu}\partial_{\sigma}M^{\sigma\mu}_{J}$, where $u^{\mu}$ is the fluid four-velocity. Which of these is the correct thermodynamic density? This issue does not arise in ordinary hydrodynamics because all fluid variables are constant in equilibrium, leading to the vanishing of such improvements. Whereas, fluid variables admit a non-trivial spacetime profile in spin hydrodynamics. Concretely, taking $M_{J}^{\mu\nu}=f(T)\partial^{[\mu}(u^{\nu]}/T)$, where $f(T)$ is an arbitrary function of the fluid temperature $T$, the particle density shifts to $n\to n-f^{\prime}(T)a^{2}$, where $a^{2}$ is the squared fluid acceleration. This ambiguity means that $n$ may violate standard thermodynamic relations in a generic pseudo-gauge.
@@
.—Conserved currents, and hence the constitutive relations of hydrodynamics, are only defined up to improvements terms that leave eq.˜1 invariant. In spin hydrodynamics, this is known as the pseudo-gauge ambiguity and takes the form Halbwachs [1960], Hehl [1976], Becattini and Tinti [2011], Speranza and Weickgenannt [2021]
@@
Pseudo-gauge ambiguity implies that the spin current has no independent physical information of its own and can always be set to zero with a suitable choice of $M_{T}$. Therefore, it is helpful to define the Belinfante tensor Belinfante [1940]
@@
We may identify the entropy density $s$, spin density $\rho^{\mu\nu}$, particle density $n$, and energy density $\epsilon$. As we are expressing the variation $\mathrm{d}p$ in terms of a tensor $\mathrm{d}\mu_{\mu\nu}$, Lorentz symmetry requires an additional term $\pi_{\mu}\mathrm{d}u^{\mu}$, such that $\rho^{[\mu}{}_{\!\rho}\mu^{\nu]\rho}=u^{[\mu}\pi^{\nu]}$ and $\pi_{\mu}u^{\mu}=0$. We may interpret $\pi_{\mu}$ as the momentum density of the fluid in its rest frame.
@@
where $X^{\mu},Y^{\mu}\in\{a^{\mu},\omega^{\mu},\ell^{\mu}\}$ and $\mathcal{X}_{XY}=\mathcal{X}_{Y\!X}$. Let us also define $\mathcal{F}=\mathcal{P}+\omega^{2}\mathcal{X}_{\omega\omega}$ for later use. Since a symmetric 2-tensor can only have 10 independent components, we can set $\mathcal{X}_{\ell\ell}=0$. Requiring that the currents are identically conserved in equilibrium gives several identities among the components, outlined in appendix˜A. The explicit components for free Dirac and scalar fields are given in appendices˜F and G.
@@
Pseudo-gauge improvements shift the components in Pseudo-gauge ambiguity of ideal spin fluids, denoted as $\mathcal{P}^{\prime}=\mathcal{P}+\delta\mathcal{P}$, etc. See appendices˜B, C and D for the explicit expressions. We want these to obey the thermodynamic form in Ideal spin fluids and Ideal spin fluids. This leads to the following constraints
@@
where $\mathscr{D}\equiv 1-(a\cdot\omega)\partial/\partial(a\cdot\omega)$. There are also constraints involving $\mathcal{E}^{\prime}$, $\mathcal{Q}_{\ell}^{\prime}$ components, but they are identically satisfied due to the conservation identities; see appendix˜E. Eq.˜21 provides linear partial differential equations (PDEs) for pseudo-gauge parameters $\gamma_{i}$’s and $\lambda_{i}$’s. Having solved these, the EoS of ideal spin fluids is given by the indefinite integral
@@
—The PDEs (21) do not admit solutions for arbitrary conserved currents in Pseudo-gauge ambiguity of ideal spin fluids. For solutions to exist, the currents must satisfy certain thermodynamic relations. In the absence of spin, we find the familiar relations: $\lim_{\mu_{\mu\nu}\to 0}\mathcal{N}-\partial\mathcal{P}/\partial\mu=0$ and $\lim_{\mu_{\mu\nu}\to 0}\mathcal{S}-\partial\mathcal{P}/\partial T=0$. Note that these do not follow from conservation. Performing an exhaustive analysis of pseudo-gauge improvements at quadratic order in spin (see appendix˜E), we find one thermodynamic relation for the Belinfante tensor and one for the particle current
@@
.—Let us illustrate our procedure using the example of free Dirac fermions Palermo et al. [2021], Palermo [2023], Ambruş and Winstanley [2014], Ambrus and Winstanley [2019], Ambrus [2020], Buzzegoli et al. [2017], Becattini and Singh [2025]. The respective currents in Pseudo-gauge ambiguity of ideal spin fluids are given in appendix˜F. The constraint equations (21) for the massless case admit a solution for pseudo-gauge parameters
@@
where $x_{n}=nm/T$, while $K_{2}$, $K_{3}$ are modified Bessel’s functions. See appendix˜F for details. We can obtain other thermodynamic quantities from eq.˜31. The quadratic pseudo-gauge invariant is given as
@@
We have repeated this exercise for free scalar fields. The answers can be found in appendix˜G.
@@ Acknowledgements.
This work partially is funded by the European Union. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. For the purpose of open access, the authors have applied a CC BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission.
Appendix A Useful identities in thermodynamic equilibrium

### v2

@@ Abstract
Conserved currents of relativistic spin fluids derived from microscopic models appear to violate local thermodynamic relations. We present a systematic analysis of pseudo-gauge improvements in ideal spin hydrodynamics and identify a family of pseudo-gauges where standard thermodynamic relations are satisfied. We quantify pseudo-gauge ambiguities in the spin equation of state and derive universal thermodynamic relations that apply to conserved currents in any pseudo-gauge. As an application, we extract the thermodynamic variables and equations of state for free Dirac fermions and scalar fields.
@@
Relativistic heavy-ion collisions produce a short-lived quark-gluon plasma (QGP) that is well described by relativistic hydrodynamics Teaney [2003], Romatschke and Romatschke [2007]. The recent observation of global $\Lambda$-hyperon polarization by the STAR Collaboration demonstrated that the QGP fireball carries substantial angular momentum and exhibits strong vortical structure Adamczyk and others [2017], STAR Collaboration [2019], ALICE Collaboration [2020, 2022]. This discovery has motivated the development of spin hydrodynamics, a theoretical framework that describes collective spin phenomena and provides novel probes of the QGP’s gradients and rotational dynamics Becattini [2011], Florkowski et al. [2018, 2019b], Montenegro et al. [2017], Hattori et al. [2019], Florkowski et al. [2019a], Becattini and Lisa [2020], Singh et al. [2021], Fukushima and Pu [2021], Hongo et al. [2021], Gallegos et al. [2021], Florkowski et al. [2022], Gallegos et al. [2023], Peng et al. [2021], Cao et al. [2022], Weickgenannt et al. [2022b, a], Biswas et al. [2023], Florkowski [2025], Huang [2025], Singh [2025], Daher et al. [2025].
@@
Hydrodynamics is a theory of conserved currents. However, these currents admit improvements—the so-called pseudo-gauge transformations—that leave the global conserved charges unchanged halbwachs1960théorie, Hehl [1976], Becattini and Tinti [2011], Speranza and Weickgenannt [2021], Becattini and Hoyos [2025]. As a result, the conserved currents in spin hydrodynamics are not unique, even in equilibrium. In particular, the constitutive relations in generic pseudo-gauges fail to satisfy the standard thermodynamic relations implied by statistical mechanics Becattini et al. [2024], Becattini and Singh [2025], Florkowski and Hontarenko [2025], Drogosz et al. [2024]. See also Florkowski and Hontarenko [2025], Drogosz et al. [2024] for generalised vector thermodynamic relations for ideal spin fluids, which we will not delve into in this work. This subtlety prevents a direct identification of thermodynamic variables of spin hydrodynamics through naive comparison with the conserved currents obtained from microscopic models or experimental measurements.
@@
For instance, the particle current admits a pseudo-gauge ambiguity $J^{\mu}\to J^{\mu}+\partial_{\sigma}M^{\sigma\mu}_{J}$, for some antisymmetric tensor $M^{\mu\nu}_{J}$. This shifts the particle density in the fluid’s rest frame as $n\to n-u_{\mu}\partial_{\sigma}M^{\sigma\mu}_{J}$, where $u^{\mu}$ is the fluid four-velocity. Which of these is the correct thermodynamic density? This issue does not arise in ordinary hydrodynamics because all fluid variables are constant in equilibrium, leading to the vanishing of such improvements. Conversely, fluid variables admit a non-trivial spacetime profile in spin hydrodynamics. Concretely, taking $M_{J}^{\mu\nu}=f(T)\partial^{[\mu}(u^{\nu]}/T)$, where $f(T)$ is an arbitrary function of the fluid temperature $T$, the particle density shifts to $n\to n-f^{\prime}(T)a^{2}$, where $a^{2}$ is the squared fluid acceleration. This ambiguity means that $n$ may violate standard thermodynamic relations in a generic pseudo-gauge.
@@
.—Conserved currents, and hence the constitutive relations of hydrodynamics, are only defined up to improvements terms that leave eq.˜1 invariant. In spin hydrodynamics, this is known as the pseudo-gauge ambiguity and takes the form halbwachs1960théorie, Hehl [1976], Becattini and Tinti [2011], Speranza and Weickgenannt [2021]
@@
Pseudo-gauge ambiguity implies that the spin current has no independent information of its own and can always be set to zero with a suitable choice of $M_{T}$. Therefore, it is helpful to define the Belinfante tensor Belinfante [1940]
@@
We may identify the entropy density $s$, spin density $\rho^{\mu\nu}$, particle density $n$, and energy density $\epsilon$. As we are expressing the variation $\mathrm{d}p$ in terms of a tensor $\mathrm{d}\mu_{\mu\nu}$, Lorentz symmetry requires an additional term $\pi_{\mu}\mathrm{d}u^{\mu}$, such that $\rho^{[\mu}{}_{\!\rho}\mu^{\nu]\rho}=u^{[\mu}\pi^{\nu]}$ and $\pi_{\mu}u^{\mu}=0$. See appendix˜A for more details. We may interpret $\pi_{\mu}$ as the momentum density of the fluid in its rest frame.
@@
where $X^{\mu},Y^{\mu}\in\{a^{\mu},\omega^{\mu},\ell^{\mu}\}$ and $\mathcal{X}_{XY}=\mathcal{X}_{Y\!X}$. Let us also define $\mathcal{F}=\mathcal{P}+\omega^{2}\mathcal{X}_{\omega\omega}$ for later use. Since a symmetric 2-tensor can only have 10 independent components, we can set $\mathcal{X}_{\ell\ell}=0$. Requiring that the currents are identically conserved in equilibrium gives several identities among the components, outlined in appendix˜B. The explicit components for free Dirac and scalar fields are given in appendices˜G and H.
@@
Pseudo-gauge improvements shift the components in Pseudo-gauge ambiguity of ideal spin fluids, denoted as $\mathcal{P}^{\prime}=\mathcal{P}+\delta\mathcal{P}$, etc. See appendices˜C, D and E for the explicit expressions. We want these to obey the thermodynamic form in Ideal spin fluids and Ideal spin fluids. This leads to the following constraints
@@
where $\mathscr{D}\equiv 1-(a\cdot\omega)\partial/\partial(a\cdot\omega)$. There are also constraints involving $\mathcal{E}^{\prime}$, $\mathcal{Q}_{\ell}^{\prime}$ components, but they are identically satisfied due to the conservation identities; see appendix˜F. Eq.˜21 provides linear partial differential equations (PDEs) for pseudo-gauge parameters $\gamma_{i}$’s and $\lambda_{i}$’s. Having solved these, the EoS of ideal spin fluids is given by the indefinite integral
@@
—The PDEs (21) do not admit solutions for arbitrary conserved currents in Pseudo-gauge ambiguity of ideal spin fluids. For solutions to exist, the currents must satisfy certain thermodynamic relations. In the absence of spin, we find the familiar relations: $\lim_{\mu_{\mu\nu}\to 0}\mathcal{N}-\partial\mathcal{P}/\partial\mu=0$ and $\lim_{\mu_{\mu\nu}\to 0}\mathcal{S}-\partial\mathcal{P}/\partial T=0$. Note that these do not follow from conservation. Performing an exhaustive analysis of pseudo-gauge improvements at quadratic order in spin (see appendix˜F), we find one thermodynamic relation for the Belinfante tensor and one for the particle current
@@
.—Let us illustrate our procedure using the example of free Dirac fermions Palermo et al. [2021], Palermo [2023], Ambruş and Winstanley [2014], Ambrus and Winstanley [2019], Ambrus [2020], Buzzegoli et al. [2017], Becattini and Singh [2025]. The respective currents in Pseudo-gauge ambiguity of ideal spin fluids are given in appendix˜G. The constraint equations (21) for the massless case admit a solution for pseudo-gauge parameters
@@
where $x_{n}=nm/T$, while $K_{2}$, $K_{3}$ are modified Bessel’s functions. See appendix˜G for details. We can obtain other thermodynamic quantities from eq.˜31. The quadratic pseudo-gauge invariant is given as
@@
We have repeated this exercise for free scalar fields. The answers can be found in appendix˜H.
@@ Acknowledgements.
This work is partially funded by the European Union. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. For the purpose of open access, the authors have applied a CC BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission.
Appendix A Comments on spin thermodynamics
In this appendix, we provide the derivation of spin thermodynamic relations in eq.˜11. Let us start with a simple example to understand the origin of the $\pi_{\mu}\mathrm{d}u^{\mu}$ term. Consider a term in pressure, $p\sim a^{2}$. It follows that
$\mathrm{d}p\sim 2a^{\mu}\mathrm{d}a_{\mu}=2a^{\mu}\mathrm{d}\mathopen{}\mathclose{{\left(\mu_{\mu\nu}u^{\nu}}}\right)=2a^{[\mu}u^{\nu]}\mathrm{d}\mu_{\mu\nu}+2a^{\mu}\mu_{\mu\nu}\mathrm{d}u^{\nu},$ (33)
which leads to
$\rho^{\mu\nu}\sim 4a^{[\mu}u^{\nu]},\qquad\pi^{\mu}\sim 2a^{\nu}\mu_{\nu\rho}\Delta^{\rho}_{\mu}.$ (34)
This illustrates the necessity of the $\pi_{\mu}\mathrm{d}u^{\mu}$ term in thermodynamics. Note that $u_{\mu}\mathrm{d}u^{\mu}=0$ due to normalisation.
More generally, the pressure $p$ is an arbitrary function of the thermodynamic scalars: $T$, $\mu$, $a^{2}$, $\omega^{2}$, and $a\cdot\omega$. We can parametrise its differential as
$\mathrm{d}p$ $=\frac{\partial p}{\partial T}\mathrm{d}T+\frac{\partial p}{\partial\mu}\mathrm{d}\mu+\frac{\partial p}{\partial a^{2}}\mathrm{d}a^{2}+\frac{\partial p}{\partial\omega^{2}}\mathrm{d}\omega^{2}+\frac{\partial p}{\partial(a\cdot\omega)}\mathrm{d}(a\cdot\omega)$
$=\frac{\partial p}{\partial T}\mathrm{d}T+\frac{\partial p}{\partial\mu}\mathrm{d}\mu+\mathopen{}\mathclose{{\left(2\frac{\partial p}{\partial a^{2}}a^{\mu}+\frac{\partial p}{\partial(a\cdot\omega)}\omega^{\mu}}}\right)\mathrm{d}a_{\mu}+\mathopen{}\mathclose{{\left(2\frac{\partial p}{\partial\omega^{2}}\omega_{\mu}+\frac{\partial p}{\partial(a\cdot\omega)}a_{\mu}}}\right)\mathrm{d}\omega^{\mu}$
$=\frac{\partial p}{\partial T}\mathrm{d}T+\frac{\partial p}{\partial\mu}\mathrm{d}\mu+\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left(2\frac{\partial p}{\partial a^{2}}a^{\mu}+\frac{\partial p}{\partial(a\cdot\omega)}\omega^{\mu}}}\right)u^{\nu}+\mathopen{}\mathclose{{\left(2\frac{\partial p}{\partial\omega^{2}}\omega_{\rho}+\frac{\partial p}{\partial(a\cdot\omega)}a_{\rho}}}\right)\frac{1}{2}\epsilon^{\mu\nu\sigma\rho}u_{\sigma}}}\right]\mathrm{d}\mu_{\mu\nu}$
$\qquad+\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left(2\frac{\partial p}{\partial a^{2}}a^{\nu}+\frac{\partial p}{\partial(a\cdot\omega)}\omega^{\nu}}}\right)\mu_{\nu\lambda}+\mathopen{}\mathclose{{\left(2\frac{\partial p}{\partial\omega^{2}}\omega^{\nu}+\frac{\partial p}{\partial(a\cdot\omega)}a^{\nu}}}\right)\frac{1}{2}\epsilon_{\lambda\nu\rho\sigma}\mu^{\rho\sigma}}}\right]\Delta^{\lambda}_{\mu}\mathrm{d}u^{\mu}.$ (35)
Further using the identities
$a^{\nu}\mu_{\nu\rho}\Delta^{\rho}_{\mu}=-\ell_{\mu},\qquad\omega^{\nu}\mu_{\nu\rho}\Delta^{\rho}_{\mu}=0,\qquad\frac{1}{2}\omega^{\nu}\epsilon_{\lambda\nu\rho\sigma}\mu^{\rho\sigma}\Delta^{\lambda}_{\mu}=-\ell_{\mu},\qquad\frac{1}{2}a^{\nu}\epsilon_{\lambda\nu\rho\sigma}\mu^{\rho\sigma}\Delta^{\lambda}_{\mu}=0,$ (36)
this becomes
$\mathrm{d}p$ $=\frac{\partial p}{\partial T}\mathrm{d}T+\frac{\partial p}{\partial\mu}\mathrm{d}\mu+\frac{1}{2}\Big(2\mathopen{}\mathclose{{\left(\chi_{aa}a^{\mu}+\chi_{a\omega}\omega^{\mu}}}\right)u^{\nu}+\mathopen{}\mathclose{{\left(\chi_{\omega\omega}\omega_{\rho}+\chi_{a\omega}a_{\rho}}}\right)\epsilon^{\mu\nu\sigma\rho}u_{\sigma}\Big)\mathrm{d}\mu_{\mu\nu}-\mathopen{}\mathclose{{\left(\chi_{aa}+\chi_{\omega\omega}}}\right)\ell_{\mu}\mathrm{d}u^{\mu},$ (37)
where we have used
$s=\frac{\partial p}{\partial T},\qquad n=\frac{\partial p}{\partial\mu},\qquad\chi_{aa}=2\frac{\partial p}{\partial a^{2}},\qquad\chi_{\omega\omega}=2\frac{\partial p}{\partial\omega^{2}},\qquad\chi_{a\omega}=\frac{\partial p}{\partial(a\cdot\omega)}.$ (38)
Eq.˜37 is precisely the thermodynamic relation in eq.˜11, with $\rho^{\mu\nu}$ and $\pi_{\mu}$ given in Ideal spin fluids.
We can also argue the form of the thermodynamic relation in eq.˜11 via Lorentz symmetry. A Lorentz transformation $\Lambda_{\mu\nu}$ (satisfying $\Lambda^{\mu}_{\penalty 10000\ \rho}\Lambda_{\mu}{}^{\nu}=\delta_{\rho}^{\nu}$) acts on spin density and chemical potential as
$\rho^{\mu\nu}\to\Lambda^{\mu}_{\penalty 10000\ \rho}\Lambda^{\nu}_{\penalty 10000\ \sigma}\rho^{\rho\sigma},\qquad\mu_{\mu\nu}\to\Lambda_{\mu}^{\penalty 10000\ \rho}\Lambda_{\nu}^{\penalty 10000\ \sigma}\mu_{\rho\sigma}.$ (39)
Note that $p$ is a scalar and hence is invariant under Lorentz transformations. This means that the differential $\mathrm{d}p$ should not depend on $\mathrm{d}\Lambda_{\mu\nu}$ after a Lorentz transformation. On the right-hand side, $s\mathrm{d}T$ and $n\mathrm{d}\mu$ are indeed invariant, but the spin term transforms as
$\frac{1}{2}\rho^{\mu\nu}\mathrm{d}\mu_{\mu\nu}\to\frac{1}{2}\rho^{\mu\nu}\mathrm{d}\mu_{\mu\nu}+\rho_{\mu}{}^{\sigma}\mu_{\nu\sigma}\mathopen{}\mathclose{{\left(\Lambda^{\rho\mu}\mathrm{d}\Lambda_{\rho}^{\penalty 10000\ \nu}}}\right).$ (40)
Note that the differential in the parenthesis is antisymmetric in $(\mu\leftrightarrow\nu)$. Therefore, the additional term does not vanish for $\rho_{[\mu}{}^{\sigma}\mu_{\nu]\sigma}\neq 0$. We can counter this with the $\pi_{\mu}\mathrm{d}u^{\mu}$ term. Using
$u^{\mu}\to\Lambda^{\mu}_{\penalty 10000\ \rho}u^{\rho},\qquad\pi_{\mu}\to\Lambda_{\mu}^{\penalty 10000\ \rho}\pi_{\rho},$ (41)
we find
$\pi_{\mu}\mathrm{d}u^{\mu}\to\pi_{\mu}\mathrm{d}u^{\mu}-u_{\mu}\pi_{\nu}\mathopen{}\mathclose{{\left(\Lambda^{\rho\mu}\mathrm{d}\Lambda_{\rho}^{\penalty 10000\ \nu}}}\right).$ (42)
Therefore, the right-hand side of the thermodynamic relation in eq.˜11 is also invariant under Lorentz transformations, provided that we take
$u_{[\mu}\pi_{\nu]}=\rho_{[\mu}{}^{\sigma}\mu_{\nu]\sigma}.$ (43)
One may check that this rule is identically true for $\rho^{\mu\nu}$ and $\pi_{\mu}$ given in Ideal spin fluids.
As an aside, note that we can also write the same thermodynamic relations in terms of the equilibrium thermodynamic variables: $T_{0}=1/\sqrt{-\beta_{0}^{2}}$, $u_{0}^{\mu}=T_{0}\beta_{0}^{\mu}$, $\mu_{0}=T_{0}\nu_{0}$, and $\mu_{0}^{\mu\nu}=T_{0}\nu_{0}^{\mu\nu}$. To wit
$\mathrm{d}p$ $=\hat{s}\mathrm{d}T_{0}+\frac{1}{2}\hat{\rho}^{\mu\nu}\mathrm{d}\mu_{0\mu\nu}+\hat{\pi}_{\mu}\mathrm{d}u^{\mu}_{0}+\hat{n}\mathrm{d}\mu_{0},$
$\hat{\epsilon}$ $=T_{0}\hat{s}+\frac{1}{2}\mu_{0\mu\nu}\hat{\rho}^{\mu\nu}+\mu_{0}\hat{n}-p,$ (44)
where we have defined the thermodynamic densities in the inertial frame of $u^{\mu}_{0}$, given as
$\hat{\epsilon}$ $=\epsilon+\mu_{0\mu\nu}x^{\nu}\Gamma\Big(\pi^{\mu}+(\epsilon+p)u^{\mu}\Big),$
$\hat{\rho}^{\mu\nu}$ $=\Gamma\mathopen{}\mathclose{{\left(\rho^{\mu\nu}+2\pi^{[\mu}x^{\nu]}+2(\epsilon+p)u^{[\mu}x^{\nu]}}}\right),$
$\hat{\pi}_{\mu}$ $=\Gamma\Big(\pi^{\nu}+\mathopen{}\mathclose{{\left(\epsilon+p}}\right)u^{\nu}\Big)\Delta_{0\mu\nu},$
$\hat{n}$ $=\Gamma n,\qquad\hat{s}=\Gamma s,$ (45)
with the Lorentz factor
$\Gamma=\frac{T}{T_{0}}=\sqrt{\frac{1}{1-2u_{0}^{\mu}\mu_{0\mu\nu}x^{\nu}-\mu_{0}^{\mu\rho}\mu_{0\mu\nu}x_{\rho}x^{\nu}}}.$ (46)
Note that the original unhatted thermodynamic densities are defined in the local rest frame of the fluid velocity $u^{\mu}$, which is non-inertial for spinning/boosted states.
Appendix B Useful identities in thermodynamic equilibrium